OpenStudy (anonymous):

Identify the 12th term of a geometric sequence where a1 = 8 and a6 = -8,192

OpenStudy (anonymous):

the geometric sequence would be as follows: a, ar, ar^2 ,ar^3 , ar^4, ar^5 and so on. so a=8 as shown by a1. a6 = ar^5 , so -8192 = 8 * r^5 , hence you can find r which is -4 As you can see in the sequence above, for the nth term of the sequence, it would be ar^n-1 So, for the 12 term ----> ar^(11-1) = ar^10 , hence it would be 8 * (-4)^10 which would give you 8388608

OpenStudy (anonymous):

no thats not it here are the different choices

OpenStudy (anonymous):

a. 134,217,728 b. 33,554,432 c. -33,554,432 d. -134,217,728

OpenStudy (anonymous):

sorry i messed up in the numbers : it would be ar^11 which is -33554432 :)

OpenStudy (anonymous):

thanks